Please show work!

Records at Hal’s Accounting Services show the following costs for year 1:

Direct materials and supplies $ 47,000
Employee costs 2,900,000
Total overhead 1,290,000

Production was 35,000 billable hours. Fixed overhead was $780,000.

Assuming no change in billable hours in year 2, direct materials and supplies costs are expected to increase by 10 percent. Direct labor costs are expected to increase by 15 percent. Variable overhead per billable hour is expected to remain the same, but fixed overhead is expected to increase by 5 percent.


Required:

a. Year 2 production is expected to be 28,000 billable hours. What are the estimated direct materials, direct labor, variable overhead, and fixed overhead costs for year 2? (Do not round intermediate computations.)

b. Determine the total costs per billable hour for year 1 and year 2. (Round your answers to 2 decimal places.)
Answer #1

a) Year 2 Production Cost

Direct material (47000/35000*28000)*1.1 41360
Direct labor (2900000/35000*28000)*1.15 2668000
Variable overhead (1290000/35000*28000) 1032000
Fixed overhead (780000*1.05) 819000
Total production Cost 4560360

b) Total Cost per billable hours :

Year 2 = 4560360/28000 = 162.87 per billable hour

Year 1 = 5017000/35000 = 143.34 per billable hour